solve the following systems of equations.
x+2y=8 (1)
x=7-2y (2)
a.the solution is??
b. there is no solution

Answer by stanbon(75887)

solve the following systems of equations.
x+2y=8 (1)
x=7-2y (2)
---
x = -2y+8
x = -2y+7
----
Since the slopes are the same and the intercepts
are different, the lines are parallel.
----
Ans: b
=============
Cheers,
Stan H.
================
a.the solution is??
b. there is no solution